OpenStudy (anonymous):

Forty is 25% of what number?

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

You need to set up a proportion: x/40 = 25/100 Now cross-multiply to solve for "x" 40*25= 1000 1000/100 = is your answer

OpenStudy (anonymous):

i was thinking 10

OpenStudy (anonymous):

there you go :)

OpenStudy (anonymous):

i used this http://www.rapidtables.com/calc/math/Percentage_Calculator.htm

OpenStudy (anonymous):

yeah good job

OpenStudy (anonymous):

i wasnt to sure so thx :D

OpenStudy (anonymous):

mhmm :D

OpenStudy (anonymous):

Hi, The answer of your question is 10. http://www.acalculator.com/percentage-calculator.html